The sum of two numbers is 112. The second is 7 more than 4 times the first. What are the two numbers?
Answer by tutorcecilia(2152)

Sum=to add
is= equals (=)
two numbers = x + y
7 more = add 7
4 times = multiply the first number by 4
Total = 112
..
So, x+y=112
The first number = x
The second number = y = 4x+7
.
x+y=112
x+(4x+7)=112
x+4x+7=112 [combine like-terms and solve for the x-term]
5x+7=112
5x+7-7=112-7
5x=105
5x/5=105/5
x=21
.
To find the second number (y), plug (x=21) back into the original formula and solvefor the y-term:
x+y=112
21+y=112
21-21+y=112-21
y=91
.
checking:
x+y=112
21+91-112
112=112 [checks out}